Assume that a random variable is normally distributed with a mean of 1,200 and a variance of 360. Complete parts a through c below. What is the probability that a randomly selected value will be greater than 1, 253​?
Answer #1

Solution:
Given: a random variable is normally distributed with a mean of 1,200 and a variance of 360.

Thus X ~ Normal (\mu =1200 \: , \: \sigma = 360)

Find: the probability that a randomly selected value will be greater than 1, 253

P( X> 1253) = .............?

Find z score for x = 1253

z=\frac{x-\mu }{\sigma }

z=\frac{1253-1200 }{360}=\frac{53 }{360}=0.15

Thus we get:

P( X> 1253) = P( Z > 0.15)

P( X> 1253) = 1 - P( Z < 0.15)

Look in z table for z = 0.1 and 0.05 and find corresponding area.

.00 .01 .02 .03 .04 .05 .06 .07 .08 .09 Z 5 99 5359 0.0 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9 .5040 .5080 5120 .5160 5239 5279

P( Z< 0.15) = 0.5596

Thus

P( X> 1253) = 1 - P( Z < 0.15)

P( X> 1253) = 1 - 0.5596

P( X> 1253) = 0.4404